# Preserving curvature lower bounds when Ricci flowing non-smooth initial data

### v1

@@ 5. Weakly differentiable Riemannian metrics and weak curvature lower bounds
We saw in Section 3, that one can transform the Ricci flow, using the harmonic heat flow, to the Ricci-DeTurck flow, which is a parabolic partial differential system of equations. One can then examine the solution using methods from the theory of parabolic partial differential systems of equations. This was carried out in [Sim01] (see also [Sim02]) for $C^{0}$ metrics and metrics which are $1+\varepsilon(n)$-fair to a smooth background metric $h$, with bounded curvature, and in [KL12, KL15] for similar settings. In the paper of Lamm/Simon, [LS23], the case that the initial metric $g(0)$ is $W^{2,2}$ and $a$-fair to a smooth complete, metric $h$ with bounded curvature, for an arbitrary $a>0$ was considered in the case $n=4$. These methods were used, generalised and improved to the case of arbitrary dimension $n,$ assuming $g(0)$ is in $W^{1,n}$ and $a$-fair to a smooth, complete, bounded curvature metric $h$, for an arbitrary $a>0$, by Chu/Lee, [CL22]. In the paper of Jiang/Sheng/Zhang, [JSZ23], the authors considered the case that $g(0)\in W^{1,p}(M)$ with $M$ closed, smooth, and $p>n.$ In the $W^{1,p}$ setting with $p>n,$ the Sobolev-embedding theorems show one that in fact $g(0)\in C^{0}(M),$ and so the paper of Simon [Sim01] (or Koch/Lamm [KL12], [KL15]) can be used to prove existence of a flow, where $g(t)\to g(0)$ uniformly locally as $t\searrow 0$.
@@
$\int_{M}|\nabla^{2}g(0)|^{2}+|\nabla g(0)|^{2}d{\rm vol}_{h}<\infty$
@@
In the first part of the paper, one considers only the Ricci-DeTurck flow. As a first step, one notices by writing things in local Riemannian coordinates, that a local Sobolev inequality holds, and that for any $\varepsilon>0$, that after scaling the metrics $g(0)$ and $h$ by the same constant (once) (we denote the scalings of the metrics $g(0)$ and $h$ also by $g(0)$ and $h$) we may assume that
$E(x):=\int_{B_{1}(x)}|\nabla^{2}g(0)|^{2}+|\nabla g(0)|^{4}d{\rm vol}_{h}\leq\varepsilon\text{ for all }\ x\in M$ (5.1)
@@
$\int_{B_{1}(x)}|\nabla^{2}g(t)|^{2}+|\nabla g(t)|^{4}d{\rm vol}_{h}\leq\sigma(b,\varepsilon),$ (5.6)
where $\sigma(b,\varepsilon)\to 0$ as $\varepsilon\to 0$ holds under Ricci flow for $t\leq T(n,b)$ as long as $\frac{1}{b}h\leq g(t)\leq bh$ remains true, if $\int_{B_{1}(x)}|\nabla^{2}g(0)|^{2}+|\nabla g(0)|^{4}d{\rm vol}_{h}\leq\varepsilon$ holds at time zero.
An important part of the argument is to show that indeed $\frac{1}{b}h\leq g(t)\leq bh$ remains true for some $b=b(a)$ if $t\leq T(a)$ if the initial $\varepsilon$ is chosen small enough. Assuming this first fails at time $t_{0}$ then one scales everything once again by $\frac{1}{t_{0}}$, so that one may assume this first fails at time $t_{0}=1.$ Standard parabolic arguments show that we must have
@@
where $c(b(a),\varepsilon)\to 0$ as $\varepsilon\to 0,$ and the second inequality is shown by integrating $\frac{\partial}{\partial t}\int_{B_{h}(x,2)}|g(1)-h|^{2}$ from $t=0$ to $t=1$. But this tells us $|g(1)-h|^{2}\leq c_{2}a^{2}$ for $\varepsilon$ small enough, which leads to a contradiction if $b=b(a)\geq c_{3}a$, and $\varepsilon=\varepsilon(a)$ is chosen small enough.
@@
For any $1<a<\infty$ there exists a constant $\varepsilon_{1}=\varepsilon_{1}(a)>0$ with the following properties. Let $(M^{4},h)$ be a smooth four dimensional Riemannian manifold which satisfies (5.3). Let $g(0)$ be a $W_{loc}^{2,2}\cap L^{\infty}$ Riemannian metric, not necessarily smooth, which satisfies
@@
$\ \ \int_{B_{2}(x)}(|{{}^{h}\nabla}g(0)|^{2}+|{{}^{h}\nabla}^{2}g(0)|^{2})dh\leq\varepsilon\ \ \mbox{ \rm for all }\ \ x\in M,$ (b)
where $\varepsilon\leq\varepsilon_{1}.$ Then there exists a constant $T=T(a,\varepsilon)>0$ and a smooth solution $(g(t))_{t\in(0,T]}$ to (3.1) such that
@@
$\ \ \ \ \int_{B_{1}(x)}(|{{}^{h}\nabla}g(\cdot,t)|^{2}+|{{}^{h}\nabla}^{2}g(\cdot,t)|^{2})dh\leq 2\varepsilon$ (${\rm b}_{t}$)
$\ \ \ \ |{{}^{h}\nabla}^{j}g(\cdot,t)|^{2}\leq\frac{c_{j}(h,a,\varepsilon)}{t^{j}}$ (${\rm c}_{t}$)
for all $x\in M,\ t\in[0,T],$ where $c_{j}(h,\varepsilon,a)\to 0$ as $\varepsilon\to 0,$ and
@@
Let $1<a<\infty$ , $M=M^{4}$ be a four dimensional manifold, and $g(0)$ and $h$ satisfy the assumptions (5.3), (a) and (b), with $\varepsilon\leq\varepsilon_{1}$ where $\varepsilon_{1}=\varepsilon_{1}(a)>0$ is the constant coming from Theorem 5.1, and let
@@
    The limit maps $\Phi(0):M\to M,$ $\Phi(0):=\lim_{t\searrow 0}\Phi(t),$ and $W(0):M\to M,$ $W(0):=\lim_{t\searrow 0}W(t),$ from (i) are also obtained in the $W_{loc}^{1,p}$ sense for $p\in[1,\infty)$. Furthermore, for any smooth coordinates $\varphi:U\to\mathbb{R}^{n}$, and $\psi:V\to\mathbb{R}^{n}$ with $W(0)(V)\subset\subset U,$ the functions $(\ell_{0})_{ij}\circ W(0):V\to\mathbb{R}$ are in $L_{loc}^{p}$ for all $p\in[1,\infty)$ and $(g(0))_{\alpha\beta}:V\to\mathbb{R}$ and $(\ell_{0})_{ij}:U\to\mathbb{R}$ are related by the identity
@@
    The metric $d_{0}$ satisfies $d_{0}(x,y):=\liminf_{\varepsilon\searrow 0}\inf_{\gamma\in C_{\varepsilon,x,y}}L_{g(0)}(\gamma),$ where $C_{\varepsilon,x,y}$ is the space of $\varepsilon$-approximative Lebesgue curves between $x$ and $y$ with respect to $g(0)$: This space is defined/examined in Definition 8.2/Section 8 of [LS23].
@@
In the paper of [JSZ23], the argument is quite subtle and they show that the Scalar curvature remains bounded from below in the distributional sense, if $g(0)\in W^{1,p}$. One of the interesting features there, is that one needs to evolve the test functions $\varphi$ used when testing whether ${\rm R}(g(t))-k\geq 0$ in the distributional definition of ${\rm R}(g(t))\geq k.$
@@
Burkhardt-Guim [BG24] considers the case that the singular set $S$ has dimension less than $n-2-\alpha,$ for some $\alpha>0$ and the metric $g_{0}$ is defined on $\mathbb{R}^{n}$ and is $L^{\infty}$ close to the standard metric, in the sense that $|g_{0}-\delta_{0}|_{\delta_{0}}\leq\varepsilon$ for an appropriately chosen $\varepsilon$. If the scalar curvature is non-negative away from the singular set, then Theorem 1.1 of [BG24] shows, using the Ricci flow and Ricci DeTurck flow, that there are smooth metrics $g_{t},t\in(0,1)$ such that $g_{t}\to g$ in $C^{\infty}(\mathbb{R}^{n}\backslash S)$ as $t\to 0,$ with ${\rm R}(g_{t})\geq 0$.
@@ 7. Open problems
    Is it possible to prove versions of the theorems of Simon [Sim12], Bamler/Cabezas-Rivas/Wilking [BCRW19], in the setting of Lamm/Simon, Theorem 5.1 : If $\mathcal{R}+\varepsilon I\in\mathcal{C}$ for some weak metric $g(0)\in W_{glob}^{2,\frac{n}{2}}$ and $h$ is $a$-fair to $g(0)$, then is $\mathcal{R}(g(t))+\varepsilon CI\in\mathcal{C}$ for $t\in[0,S)$ for some uniform $S>0$ depending on $h$ for the Lamm/Simon, [LS23], solution ? For general $n\neq 4$ it would first be necessary to solve (b), before considering this question.
@@
    With respect to (i): in the paper [CM23] the authors showed that if one assumes a collapsing of the type $\frac{{\rm vol}(B_{g(0)}(x,r))}{r^{3}}\geq f(r)$ for a particular function $f:\mathbb{R}^{+}\to\mathbb{R}^{+}$ where $\lim_{r\to\infty}f(r)=0,$ then there is a complete solution to Ricci flow for a short time: See [CM23] for details.

### v2

@@ 5. Weakly differentiable Riemannian metrics and weak curvature lower bounds
We saw in Section 3, that one can transform the Ricci flow, using the harmonic heat flow, to the Ricci-DeTurck flow, which is a parabolic partial differential system of equations. One can then examine the solution using methods from the theory of parabolic partial differential systems of equations. This was carried out in [Sim01] (see also [Sim02]) for $C^{0}$ metrics and metrics which are $1+\epsilon(n)$-fair to a smooth background metric $h$, with bounded curvature, and in [KL12, KL15] for similar settings. In the paper of Lamm/Simon, [LS23], the case that the initial metric $g(0)$ is $W^{2,2}$ and $a$-fair to a smooth complete, metric $h$ with bounded curvature, for an arbitrary $a>0$ was considered in the case $n=4$. These methods were used, generalised and improved to the case of arbitrary dimension $n,$ assuming $g(0)$ is in $W^{1,n}$ and $a$-fair to a smooth, complete, bounded curvature metric $h$, for an arbitrary $a>0$, by Chu/Lee, [CL22]. In the paper of Jiang/Sheng/Zhang, [JSZ23], the authors considered the case that $g(0)\in W^{1,p}(M)$ with $M$ closed, smooth, and $p>n.$ In the $W^{1,p}$ setting with $p>n,$ the Sobolev-embedding theorems show one that in fact $g(0)\in C^{0}(M),$ and so the paper of Simon [Sim01] (or Koch/Lamm [KL12], [KL15]) can be used to prove existence of a flow, where $g(t)\to g(0)$ uniformly locally as $t\searrow 0$.
@@
$\int_{M}|\nabla^{2}g(0)|^{2}+|\nabla g(0)|^{2}dV_{h}<\infty$
@@
In the first part of the paper, one considers only the Ricci-DeTurck flow. As a first step, one notices by writing things in local Riemannian coordinates, that a local Sobolev inequality holds, and that for any $\epsilon>0$, that after scaling the metrics $g(0)$ and $h$ by the same constant (once) (we denote the scalings of the metrics $g(0)$ and $h$ also by $g(0)$ and $h$) we may assume that
$E(x):=\int_{B_{1}(x)}|\nabla^{2}g(0)|^{2}+|\nabla g(0)|^{4}dV_{h}\leq\epsilon\text{ for all }\ x\in M$ (5.1)
@@
$\int_{B_{1}(x)}|\nabla^{2}g(t)|^{2}+|\nabla g(t)|^{4}dV_{h}\leq\sigma(b,\epsilon),$ (5.6)
where $\sigma(b,\epsilon)\to 0$ as $\epsilon\to 0$ holds under Ricci flow for $t\leq T(n,b)$ as long as $\frac{1}{b}h\leq g(t)\leq bh$ remains true, if $\int_{B_{1}(x)}|\nabla^{2}g(0)|^{2}+|\nabla g(0)|^{4}dV_{h}\leq\epsilon$ holds at time zero.
An important part of the argument is to show that indeed $\frac{1}{b}h\leq g(t)\leq bh$ remains true for some $b=b(a)$ if $t\leq T(a)$ if the initial $\epsilon$ is chosen small enough. Assuming this first fails at time $t_{0}$ then one scales everything once again by $\frac{1}{t_{0}}$, so that one may assume this first fails at time $t_{0}=1.$ Standard parabolic arguments show that we must have
@@
where $c(b(a),\epsilon)\to 0$ as $\epsilon\to 0,$ and the second inequality is shown by integrating $\frac{\partial}{\partial t}\int_{B_{h}(x,2)}|g(1)-h|^{2}$ from $t=0$ to $t=1$. But this tells us $|g(1)-h|^{2}\leq c_{2}a^{2}$ for $\epsilon$ small enough, which leads to a contradiction if $b=b(a)\geq c_{3}a$, and $\epsilon=\epsilon(a)$ is chosen small enough.
@@
For any $1<a<\infty$ there exists a constant $\epsilon_{1}=\epsilon_{1}(a)>0$ with the following properties. Let $(M^{4},h)$ be a smooth four dimensional Riemannian manifold which satisfies (5.3). Let $g(0)$ be a $W_{loc}^{2,2}\cap L^{\infty}$ Riemannian metric, not necessarily smooth, which satisfies
@@
$\ \ \int_{B_{2}(x)}(|{{}^{h}\nabla}g(0)|^{2}+|{{}^{h}\nabla}^{2}g(0)|^{2})dh\leq\epsilon\ \ \mbox{ \rm for all }\ \ x\in M,$ (b)
where $\epsilon\leq\epsilon_{1}.$ Then there exists a constant $T=T(a,\epsilon)>0$ and a smooth solution $(g(t))_{t\in(0,T]}$ to (3.1) such that
@@
$\ \ \ \ \int_{B_{1}(x)}(|{{}^{h}\nabla}g(\cdot,t)|^{2}+|{{}^{h}\nabla}^{2}g(\cdot,t)|^{2})dh\leq 2\epsilon$ (${\rm b}_{t}$)
$\ \ \ \ |{{}^{h}\nabla}^{j}g(\cdot,t)|^{2}\leq\frac{c_{j}(h,a,\epsilon)}{t^{j}}$ (${\rm c}_{t}$)
for all $x\in M,\ t\in[0,T],$ where $c_{j}(h,\epsilon,a)\to 0$ as $\epsilon\to 0,$ and
@@
Let $1<a<\infty$ , $M=M^{4}$ be a four dimensional manifold, and $g(0)$ and $h$ satisfy the assumptions (5.3), (a) and (b), with $\epsilon\leq\epsilon_{1}$ where $\epsilon_{1}=\epsilon_{1}(a)>0$ is the constant coming from Theorem 5.1, and let
@@
    The limit maps $\Phi(0):M\to M,$ $\Phi(0):=\lim_{t\searrow 0}\Phi(t),$ and $W(0):M\to M,$ $W(0):=\lim_{t\searrow 0}W(t),$ from (i) are also obtained in the $W_{loc}^{1,p}$ sense for $p\in[1,\infty)$. Furthermore, for any smooth coordinates $\phi:U\to\mathbb{R}^{n}$, and $\psi:V\to\mathbb{R}^{n}$ with $W(0)(V)\subset\subset U,$ the functions $(\ell_{0})_{ij}\circ W(0):V\to\mathbb{R}$ are in $L_{loc}^{p}$ for all $p\in[1,\infty)$ and $(g(0))_{\alpha\beta}:V\to\mathbb{R}$ and $(\ell_{0})_{ij}:U\to\mathbb{R}$ are related by the identity
@@
    The metric $d_{0}$ satisfies $d_{0}(x,y):=\liminf_{\epsilon\searrow 0}\inf_{\gamma\in C_{\epsilon,x,y}}L_{g(0)}(\gamma),$ where $C_{\epsilon,x,y}$ is the space of $\epsilon$-approximative Lebesgue curves between $x$ and $y$ with respect to $g(0)$: This space is defined/examined in Definition 8.2/Section 8 of [LS23].
@@
In the paper of [JSZ23], the argument is quite subtle and they show that the Scalar curvature remains bounded from below in the distributional sense, if $g(0)\in W^{1,p}$. One of the interesting features there, is that one needs to evolve the test functions $\phi$ used when testing whether ${\rm R}(g(t))-k\geq 0$ in the distributional definition of ${\rm R}(g(t))\geq k.$
@@
Burkhardt-Guim [BG24] considers the case that the singular set $S$ has dimension less than $n-2-\alpha,$ for some $\alpha>0$ and the metric $g_{0}$ is defined on $\mathbb{R}^{n}$ and is $L^{\infty}$ close to the standard metric, in the sense that $|g_{0}-\delta_{0}|_{\delta_{0}}\leq\epsilon$ for an appropriately chosen $\epsilon$. If the scalar curvature is non-negative away from the singular set, then Theorem 1.1 of [BG24] shows, using the Ricci flow and Ricci DeTurck flow, that there are smooth metrics $g_{t},t\in(0,1)$ such that $g_{t}\to g$ in $C^{\infty}(\mathbb{R}^{n}\backslash S)$ as $t\to 0,$ with ${\rm R}(g_{t})\geq 0$.
@@ 7. Open problems
    Is it possible to prove versions of the theorems of Simon [Sim12], Bamler/Cabezas-Rivas/Wilking [BCRW19], in the setting of Lamm/Simon, Theorem 5.1 : If $\mathcal{R}+\epsilon I\in\mathcal{C}$ for some weak metric $g(0)\in W_{glob}^{2,\frac{n}{2}}$ and $h$ is $a$-fair to $g(0)$, then is $\mathcal{R}(g(t))+\epsilon CI\in\mathcal{C}$ for $t\in[0,S)$ for some uniform $S>0$ depending on $h$ for the Lamm/Simon, [LS23], solution ? For general $n\neq 4$ it would first be necessary to solve (b), before considering this question.
@@
    With respect to (i): in the paper [CM23] the authors showed that if one assumes a collapsing of the type $\frac{V(B_{g(0)}(x,r))}{r^{3}}\geq f(r)$ for a particular function $f:\mathbb{R}^{+}\to\mathbb{R}^{+}$ where $\lim_{r\to\infty}f(r)=0,$ then there is a complete solution to Ricci flow for a short time: See [CM23] for details.